# Speedups in nonequilibrium thermal relaxation: Mpemba and related effects

### v1

@@ 3.2 Relaxation through a second-order phase transition
As far as we know, a Mpemba effect through a second-order phase transition has never been experimentally observed, despite a theoretical argument for its existence [57]. Identifying the phase transition time is a main challenge in addressing a Mpemba effect through a second-order phase transition. The two mechanisms suggested for a first-order phase transition bypassed the need to exactly define the phase transition time to observe a Mpemba effect: The “finite bath” mechanism suggested in [127] assumes that the hot and cold temperature reach essentially the same metastable state and that it is a property of the thermal bath (its temperature) that differentiates the evolution from the metastable state to the final equilibrium state. This makes it unnecessary to exactly define the phase transition time (which indeed diverges in the thermodynamic limit) because it is clear that a rare fluctuation that takes the system from the metastable state to the final equilibrium state is more probable – and therefore, its average time is smaller – when the system is coupled to higher temperature bath. In the “different basin of attraction” mechanism suggested in [58], the exact definition of the phase transition time is again not essential, since it is clear that the hot initial condition relaxes directly to the final equilibrium, whereas the cold initial condition relaxes first to a metastable state and thus takes longer (in the thermodynamic limit, an infinite time) to reach the equilibrium state, no matter how the phase transition time is defined.
@@
In contrast to the two mechanisms for a Mpemba effect through a first-order phase transition discussed above, the definition of the phase transition time plays a key role in the Mpemba effect through a second-order phase transition. These transitions are usually not associated with a metastable state, and thus, the transition is not expected to be dominated by fluctuations, as in a first-order phase transition. However, defining the exact time at which the phase transition happens is somewhat subtle since many of the equilibrium characteristics of an equilibrium second-order phase transition, e.g., divergence of the correlation length or the percolation transition, take infinite time in the thermodynamic limit. The definition of the phase transition time used in [57] is based on the following picture of a second-order phase transition: the system has an order parameter whose physical value minimizes the (Landau) free energy. At the phase transition temperature, the minimum of the free energy splits into two minima, with a maximum in between. The maximum corresponds to the instability of the system in the disordered state. The phase transition time is defined as the time at which the disordered phase becomes unstable. For systems with a single order parameter, the phase transition time must then be zero. However, as demonstrated in Fig. 19, for systems with more than a single order parameter, this definition can provide a non-zero time for the phase transition, which does not diverge in the thermodynamic limit. It can thus be used to compare the phase transition time for different initial conditions and, therefore, to define and detect a Mpemba effect.
@@
Figure 19: A system with two order parameters can have a phase transition at a finite time. The phase of the system is defined based on the value of the order parameter $x_{1}$ and its stability: $x_{1}=0$ is the disordered phase, and $x_{1}\neq 0$ is the ordered phase. However, the stability of $x_{1}=0$ can change as a function of the other order parameter, $x_{2}$, as shown in the figure: in the gray area $x_{1}=0$ is still stable, and in the yellow area it is not. Therefore, the phase transition happens when the system crosses the black line. A representative trajectory is demonstrated by the pink line. If the two initial conditions are arranged as the blue (cold) and red (hot) points in the figure, then there will be a Mpemba effect since the initially hot system crosses the phase transition line first. This figure is reproduced from [57].
@@
Let us demonstrate the phase transition time on a simple system – the mean field antiferromagnetic Ising model. In this model, each spin lives on a node. Half of the nodes belong to sublattice A, and the other half to sublattice B. The two sublattices represent the two sublattices of an antiferromagnetic system. Each spin in sublattice A is coupled antiferromagnetically to all spins in sublattice B and vice versa, but there is no coupling between the spins in the same sublattice. In addition, there is an external magnetic field coupled to all the spins in the system. We denote the number of spins up (down) in sublattice A as $N^{A}_{\uparrow}$ ($N^{A}_{\downarrow}$) and for sublattice B as $N^{B}_{\uparrow}$ ($N^{B}_{\downarrow}$). The total magnetization is denoted as $m=(N^{A}_{\uparrow}+N^{B}_{\uparrow}-N^{A}_{\downarrow}-N^{B}_{\downarrow})/N$ and the staggered magnetization, which is the magnetization difference between the two sublattices, as $m_{s}=(N^{A}_{\uparrow}-N^{B}_{\uparrow}-N^{A}_{\downarrow}+N^{B}_{\downarrow})/N$. With these, the Hamiltonian of the system is given by
@@
Figure 20: The arrows represent the dynamic of the staggered magnetization and magnetization, from  Eqs. (75) and (76), on the $(m_{s},m)$ plane. The green line is the value of the magnetization, $m$, at which the stability of the $m_{s}=0$ changes: $m_{s}=0$ is stable for smaller values of $m$ and unstable for higher values of $m$. Therefore, the phase transition time is the time $t^{*}$ at which (on average) the system crosses the green line. Two realizations (with noise) are plotted – the red line initiated from the hot initial condition and the blue initiated from the cold. (b) Zoom in around $m_{s}=0$. The figure is adapted from [57].
The mean-field, antiferromagnetic Ising model described above demonstrates that a second-order phase transition defined through the instability of the order parameter is possible at finite, non-zero time. Once the phase transition time has been defined, it is easy to identify direct and inverse Mpemba effects. However, a concrete microscopic model with dynamics that demonstrates any of these effects is not yet known. Moreover, it is not even obvious if this definition is useful beyond mean-field models. A direct Mpemba effect of this type was nevertheless demonstrated through a phenomenological Landau free energy function (shown in Fig. 19), where the temperature dependence was constructed such that a direct Mpemba effect through a second-order phase transition exists.
@@ 4 Kinetic theory: granular and molecular gases
So far, we have discussed two frameworks that are relevant to Mpemba effects: the Markovian framework and Mpemba effects through phase transitions. In this section, we present a third framework, driven granular gases, which has been used to develop Mpemba-related concepts in parallel to the Markovian framework. Because the Markovian framework discussed in Sec. 2 requires knowing all the details of the microstates and the transition rates connecting them, it is useful primarily for small or few-body systems. The thermodynamic limit can be taken in only a very few cases, such as 1D or mean-field systems. The phase-transition analysis in Sec. 3, on the other hand, relies on the dynamics of a macroscopic system away from equilibrium and thus has been applied so far only to mean-field models. The driven-granular-gas framework offers a different yet complementary approach that allows analytic study of many-body systems in the thermodynamic limit. An important difference from the Markovian systems of Sec. 2 and the phase transitions of Sec. 3 is that granular gases are inherently out of equilibrium and need the injection of energy to sustain a fluidized state, implying that stationary states are nonequilibrium steady states.
@@
The kinetic theory of gases offers a natural framework for studying these phenomena. This well-developed and established theory, in turn, offers insights into the emergence of anomalous relaxation effects. In the illustrative example of a granular gas of rough spheres [137, 138] (see below), the equilibrium temperature can be characterized by the average velocity of the spheres or, equivalently, by their angular velocity. However, when an external parameter of such a system is quenched to a different value, the relaxation times of these velocities can be drastically different, and the energy transfer between the angular and linear velocity due to collisions is highly nonlinear. This nonlinearity can generate anomalous relaxation phenomena.
Two different mechanisms for the emergence of anomalous effects in granular gases have been studied. The first leverages the coupling between the non-Gaussian part of the velocity distribution and the kinetic temperature [4, 139, 136, 140, 141, 142]. This mechanism works even with smooth spheres, where the angular velocities do not play any role [4]. The second mechanism exploits the energy exchange between linear and angular components of the kinetic energy [143, 137, 138]. Prototypical examples include several models of granular spheres, hard and viscoelastic. The latter include Maxwell models with and without external driving. Another example is rough, hard spheres with translational driving. Others include disks and also rotational driving. Finally, there are molecular gases, with ingredients such as drag force, viscosity, and even mixtures of particles. In the context of anomalous relaxations, the main differences between granular and molecular fluids is that, in the former, the Mpemba effect can emerge naturally, as granular fluids are intrinsically out of equilibrium. The dissipative interactions between material particles need an injection of energy to sustain a fluidized state. Note that even in a homogeneous cooling state of granular fluids [144], the Mpemba can be observed. In contrast, in molecular gases, the interactions and collisions are elastic (conserve energy), and the system must be driven out of equilibrium by introducing additional ingredients, namely, external nonlinear drag [139], another species [145], or immersing it in a surrounded fluid [139, 140].
@@ 4.1 The Boltzmann-Fokker-Planck equation and hydrodynamics
The kinetic theory of gases is a classical theory capable of modeling dilute gases [146]. In this theory, the system is composed of an ensemble of particles interacting through binary collisions obeying the molecular-chaos hypothesis. The Boltzmann-Fokker-Planck equation for the homogeneous states of an ensemble of hard spheres is
$\left(\partial_{t}-\nabla_{\vec{v}}\right)\left(\vec{F}(\vec{v})+\frac{\chi_{0}^{2}}{2}\nabla_{\vec{v}}\right)f(\vec{v},\vec{w};t)={\mathcal{J}[\vec{v},\vec{w}|f(t)]},$ (77)
@@
  • $f(\vec{v},\vec{w};t)$ is the one-body velocity distribution function, $\vec{v}$ and $\vec{w}$ being the translational and angular velocities, respectively;
@@
  • $\vec{F}(\vec{v})$ is a velocity-dependent force;
@@
The stochastic force results from an external stochastic driving that compensates for the energy that is lost in each collision and thereby avoids relaxation towards zero temperature. It represents the influence of a stochastic, white noise $\vec{F}^{\mathrm{wn}}$ that acts on each particle, and satisfies $\langle\vec{F}_{i}^{\mathrm{wn}}(t)\rangle=\vec{0}$ and $\langle F_{i,q}^{\mathrm{wn}}(t)F_{j,s}^{\mathrm{wn}}(t^{\prime})\rangle=m^{2}\chi_{0}^{2}\delta_{i\!j}\delta_{q,s}\delta(t-t^{\prime})$, where indices $i,j$ refer to particle numbers and the indices $q,s\in\{x,y,z\}$ to the spatial coordinates. Note that, for simplicity, we assume a stochastic force but no stochastic moment; thus, there is no diffusion term for the angular velocity. Additional rotational driving can be considered [138], but translational driving illustrates better the situation encountered in most experimental systems and physical scenarios, in which the stochastic force tries to mimic the injection of energy produced in granular gases immersed in thermal fluids or a fluidization produced by vibrations.
@@
Equation (77) describes the evolution of the single-particle velocity distribution for a collection of identical spheres at low particle density. The density is assumed constant so that the collision rate is independent of the particle position. Additionally, Eq. (77) assumes the molecular chaos hypothesis and that all collisions are binary and instantaneous. An important observation is that while this equation shares many features with the Fokker-Planck Eq. (47), there is a fundamental difference: Eq. (77) describes the evolution in time of a velocity distribution $f(\vec{v},\vec{w};t)$, regulating the probability of finding a particle at time $t$ with velocities $\vec{v}$ and $\vec{w}$, whereas Eq. (47) is an evolution equation for the probability to find the system in some specific microscopic configuration $\vec{x}$. An additional key difference between the two descriptions lies in the physical mechanisms that can enable the emergence of anomalous relaxation effects. On the one hand, in the Markovian framework, the potential $U(\vec{x})$ in Eq. (47) can be arbitrarily complicated (Fig. 6a), and the roughness of the potential is ultimately responsible for much of the anomalous effects that one observes in these setups. On the other hand, in Eq. (77), anomalous effects can be caused by a nonlinear, drag-velocity-dependent force $\vec{F}(\vec{v})$ in molecular gases. They can also be caused by a shear force and/or the inelasticity of collisions in granular gases (expressed in $\mathcal{J}[\vec{v},\vec{w}|f(t)]$) that produces a non-Gaussian distribution function that couples the kinetic temperature with higher-order moments.
@@
To keep the derivation simple, we analyze a relatively easy case, where there is no deterministic force, and the energy that is lost in each collision is compensated only by the stochastic homogeneous volume force. This simplification gives a Boltzmann-Fokker-Planck equation of the form
@@
which roughly accounts for the accumulated number of collisions per particle. In the case of thermostated systems, time can be rescaled with the stationary temperature $T_{s}$,where $T_{s}$ corresponds to the temperature the system reaches when dissipation and injection of energy are balanced; see Sec. 4.2 and 4.3. Dimensionless velocities then follow
@@
$\phi(\vec{c},\vec{j};\tau)\equiv\frac{1}{n}\left[\frac{4T_{t}(t)T_{r}(t)}{mI}\right]^{3/2}f(\vec{v},{\vec{j}};t).$ (86)
@@
Our goal is to use dimensionless variables and write evolution equations for the various temperatures defined above, which can offer a means to characterize crossings and hence anomalous relaxation effects. To achieve this goal, the exact solution to the Boltzmann-Fokker-Planck equation Eq. (78) can be formally expanded around a Maxwellian distribution, which here is a Gaussian distribution with variances $T_{t}$ and $T_{r}$ for the linear and angular velocities, respectively [149]. Explicitly, we write
$\phi_{M}(c,j)=n\pi^{-3}e^{-c^{2}-j^{2}}\sqrt{mI/4(T_{r}T_{t})}.$ (88)
@@
are dimensionless measures of noise intensity.
Given the temporal evolution of the temperature, we can define various anomalous relaxation phenomena: If, in two relaxation processes with different initial conditions but identical steady state, the corresponding curves of $\theta(\tau)$ cross, then there is a direct/inverse Mpemba effect. The specific type of the effect depends on the relaxation process – heating or cooling. The Mpemba effect has been studied using numerical, analytical techniques and two simulation methods detailed in the numerical observation; see Sec. 7.3.
@@
where the asterisks ($\vec{v}^{*},\vec{w}^{*}$) indicate precollisional velocities, namely, the velocities that after collision give as results ($\vec{v},\vec{w}$). Note that in the case of molecular gases, the collision kernel is the same, but $\alpha$ must be replaced by $1$, which corresponds to the case of elastic collisions.

### v3

@@ 3.2 Relaxation through a second-order phase transition
As far as we know, a Mpemba effect through a second-order phase transition has never been experimentally observed, despite a theoretical argument for its existence [57]. Identifying the phase transition time is a main challenge in addressing a Mpemba effect through a second-order phase transition. The two mechanisms suggested for a first-order phase transition bypassed the need to exactly define the phase transition time to observe a Mpemba effect: The “finite bath” mechanism suggested in [132] assumes that the hot and cold temperature reach essentially the same metastable state and that it is a property of the thermal bath (its temperature) that differentiates the evolution from the metastable state to the final equilibrium state. This makes it unnecessary to exactly define the phase transition time, which indeed diverges in the thermodynamic limit, because it is clear that a rare fluctuation that takes the system from the metastable state to the final equilibrium state is more probable—and therefore, its average time is smaller—when the system is coupled to higher temperature bath. In the “different basin of attraction” mechanism suggested in [58], the exact definition of the phase transition time is again not essential, since it is clear that the hot initial condition relaxes directly to the final equilibrium, whereas the cold initial condition relaxes first to a metastable state and thus takes longer (in the thermodynamic limit, an infinite time) to reach the equilibrium state, no matter how the phase transition time is defined.
@@
In contrast to the two mechanisms for a Mpemba effect through a first-order phase transition discussed above, the definition of the phase-transition time plays a key role in the Mpemba effect through a second-order phase transition. These transitions are usually not associated with a metastable state, and thus, the transition is not expected to be dominated by fluctuations, as in a first-order phase transition. However, defining the exact time at which the phase transition happens is somewhat subtle since many of the equilibrium characteristics of an equilibrium second-order phase transition, e.g., divergence of the correlation length or the percolation transition, take infinite time in the thermodynamic limit. The definition of the phase-transition time used in [57] is based on the following picture of a second-order phase transition: the system has an order parameter whose physical value minimizes the (Landau) free energy. At the phase transition temperature, the minimum of the free energy splits into two minima, with a maximum in between. The maximum corresponds to the instability of the system in the disordered state. The phase transition time is defined as the time at which the disordered phase becomes unstable. For systems with a single order parameter, the phase transition time must then be zero. However, as demonstrated in Fig. 19, for systems with more than a single order parameter, this definition can provide a non-zero time for the phase transition, which does not diverge in the thermodynamic limit. It can thus be used to compare the phase transition time for different initial conditions and, therefore, to define and detect a Mpemba effect.
@@
Figure 19: A system with two order parameters can have a phase transition at a finite time. The phase of the system is defined based on the value of the order parameter $x_{1}$ and its stability: $x_{1}=0$ is the disordered phase, and $x_{1}\neq 0$ is the ordered phase. However, the stability of $x_{1}=0$ can change as a function of the other order parameter, $x_{2}$, as shown in the figure: in the gray area $x_{1}=0$ is still stable, and in the yellow area it is not. Therefore, the phase transition happens when the system crosses the black line. A representative trajectory is demonstrated by the pink line. If the two initial conditions are arranged as the blue (cold) and red (hot) points in the figure, then there will be a Mpemba effect since the initially hot system crosses the phase transition line first. Source: Reprinted with permission from [57].
@@
Let us demonstrate the phase transition time on a simple system, the mean field antiferromagnetic Ising model. In this model, each spin lives on a node. Half of the nodes belong to sublattice A, and the other half to sublattice B. The two sublattices represent the two sublattices of an antiferromagnetic system. Each spin in sublattice A is coupled antiferromagnetically to all spins in sublattice B and vice versa, but there is no coupling between the spins in the same sublattice. In addition, there is an external magnetic field coupled to all the spins in the system. We denote the number of spins up (down) in sublattice A as $N^{A}_{\uparrow}$ ($N^{A}_{\downarrow}$) and for sublattice B as $N^{B}_{\uparrow}$ ($N^{B}_{\downarrow}$). The total magnetization is denoted as $m=(N^{A}_{\uparrow}+N^{B}_{\uparrow}-N^{A}_{\downarrow}-N^{B}_{\downarrow})/N$ and the staggered magnetization, which is the magnetization difference between the two sublattices, as $m_{s}=(N^{A}_{\uparrow}-N^{B}_{\uparrow}-N^{A}_{\downarrow}+N^{B}_{\downarrow})/N$. With these, the Hamiltonian of the system is given by
@@
Figure 20: Mpemba effect in a mean-field antiferromagnetic Ising model. (a) The arrows represent the dynamics of the staggered magnetization and magnetization, from  Eqs. (75) and (76), on the $(m_{s},m)$ plane. The green line is the value of the magnetization, $m$, at which the stability of the $m_{s}=0$ changes: $m_{s}=0$ is stable for smaller values of $m$ and unstable for higher values of $m$. Therefore, the phase transition time is the time $t^{*}$ at which (on average) the system crosses the green line. Two realizations (with noise) are plotted. The red line is initiated from the hot initial condition and the blue is initiated from the cold. (b) Zoom in around $m_{s}=0$. Source: Adapted with permission from [57].
The mean-field, antiferromagnetic Ising model described above demonstrates that a second-order phase transition defined through the instability of the order parameter is possible at finite, non-zero time. Once the phase transition time has been defined, it is easy to identify direct and inverse Mpemba effects. However, a concrete microscopic model with dynamics that demonstrates any of these effects is not yet known. Moreover, it is not even obvious if this definition is useful beyond mean-field models. A direct Mpemba effect of this type was nevertheless demonstrated through a phenomenological Landau free energy function, shown in Fig. 19, where the temperature dependence was constructed such that a direct Mpemba effect through a second-order phase transition exists.
@@ 4 Kinetic theory: granular and molecular gases
So far, we have discussed two frameworks that are relevant to Mpemba effects: the Markovian framework and Mpemba effects through phase transitions. Here we present a third framework, driven granular and molecular gases, which has been used to develop Mpemba-related concepts in parallel to the Markovian framework. Because the Markovian framework discussed in Sec. 2 requires knowledge of all details of the microstates and the transition rates connecting them, it is useful primarily for small or few-body systems. The thermodynamic limit can be taken in only a very few cases, such as 1D or mean-field systems. The phase-transition analysis in Sec. 3, on the other hand, relies on the dynamics of a macroscopic system away from equilibrium and thus has been applied so far only to mean-field models. The driven-granular-gas framework offers a different yet complementary approach that allows analytic study of many-body systems in the thermodynamic limit. An important difference from the Markovian systems of Sec. 2 and the phase transitions of Sec. 3 is that granular gases are inherently out of equilibrium and need the injection of energy to sustain a fluidized state, implying that stationary states are nonequilibrium steady states.
@@
The kinetic theory of gases offers a natural framework for studying these phenomena. This well-developed and established theory, in turn, offers insights into the emergence of anomalous relaxation effects. In the illustrative example of a granular gas of rough spheres [144, 145] (see below), the kinetic temperature is characterized by the average of translational and rotational kinetic energies of the spheres. However, when an external parameter of such a system is quenched to a different value, the relaxation times of these velocities can be drastically different, and the energy transfer between the angular and linear velocity due to collisions is highly nonlinear. This nonlinearity can generate anomalous relaxation phenomena.
Two different mechanisms for the emergence of anomalous effects in granular gases have been studied. The first leverages the coupling between the non-Gaussian part of the velocity distribution and the kinetic temperature [4, 146, 143, 147, 148, 149]. This mechanism works even with smooth spheres, where the angular velocities do not play any role [4]. The second mechanism exploits the energy exchange between linear and angular components of the kinetic energy [150, 144, 145]. Prototypical examples include several models of granular spheres, hard and viscoelastic. The latter include Maxwell models with and without external driving. Another example is rough, hard spheres with translational driving. Others include disks and also rotational driving. Finally, there are molecular gases, with ingredients such as drag force, viscosity, and even mixtures of particles. In the context of anomalous relaxations, the main difference between granular and molecular fluids is that, in the former, the Mpemba effect can emerge naturally, as granular fluids are intrinsically out of equilibrium. Namely, even when granular gases evolve freely, the Mpemba effect can appear. The dissipative interactions between material particles need an injection of energy to sustain a fluidized state. Note that even in a homogeneous cooling state of granular fluids [151], the Mpemba can be observed. In contrast, in molecular gases, the interactions and collisions are elastic (conserve energy), and the system must be driven out of equilibrium by introducing additional ingredients, namely, external nonlinear drag [146], another species [152], or immersing it in a surrounding fluid [146, 147].
@@ 4.1 The Boltzmann-Fokker-Planck equation and hydrodynamics
The kinetic theory of gases is a classical approach to modeling dilute gases [153]. In this theory, the system is composed of an ensemble of particles interacting through binary collisions obeying the molecular-chaos hypothesis. The Boltzmann-Fokker-Planck equation for the homogeneous states of an ensemble of hard spheres is
$\partial_{t}f(\vec{v},\vec{w};t)-\nabla_{\vec{v}}\cdot\left(\frac{\vec{F}(\vec{v})}{m}+\frac{\chi_{0}^{2}}{2}\nabla_{\vec{v}}\right)f(\vec{v},\vec{w};t)={\mathcal{J}[\vec{v},\vec{w}|f(t)]},$ (77)
@@
  • 1.
    $f(\vec{v},\vec{w};t)$ is the one-body velocity distribution function, with $\vec{v}$ and $\vec{w}$ the translational and angular velocities, respectively;
  • 2.
@@
  • 3.
    $\vec{F}(\vec{v})$ is a velocity-dependent force and $m$ the mass of particles;
  • 4.
@@
The stochastic force results from an external stochastic driving that compensates for the energy that is lost in each collision and thereby avoids relaxation towards zero temperature. It represents the influence of a stochastic, white noise $\vec{F}^{\mathrm{wn}}$ that acts on each particle, and it satisfies $\langle\vec{F}_{i}^{\mathrm{wn}}(t)\rangle=\vec{0}$ and $\langle F_{i,q}^{\mathrm{wn}}(t)F_{j,s}^{\mathrm{wn}}(t^{\prime})\rangle=m^{2}\chi_{0}^{2}\delta_{i\!j}\delta_{q,s}\delta(t-t^{\prime})$, where indices $i,j$ refer to particle numbers and the indices $q,s\in\{x,y,z\}$ to the spatial coordinates. Note that, for simplicity, we assume a stochastic force but no stochastic torque; thus, there is no diffusion term for the angular velocity. Additional rotational driving can be considered [145], but translational driving illustrates better the situation encountered in most experimental systems and physical scenarios, in which the stochastic force tries to mimic the injection of energy produced in granular gases immersed in thermal fluids or a fluidization produced by vibrations.
@@
Equation (77) describes the evolution of the single-particle velocity distribution for a collection of identical spheres at low particle density. The density is assumed constant, so that the collision rate is independent of the particle position. Additionally, Eq. (77) assumes the molecular chaos hypothesis and that all collisions are binary and instantaneous. An important observation is that while this equation shares many features with the Fokker-Planck Eq. (47), there is a fundamental difference: Eq. (77) describes the evolution in time of a velocity distribution $f(\vec{v},\vec{w};t)$, regulating the probability of finding a particle at time $t$ with velocities $\vec{v}$ and $\vec{w}$, whereas Eq. (47) is an evolution equation for the probability to find the system in some specific microscopic configuration $\vec{x}$. An additional key difference between the two descriptions lies in the physical mechanisms that can enable the emergence of anomalous relaxation effects. On the one hand, in the Markovian framework, the potential $U(\vec{x})$ in Eq. (47) can be arbitrarily complicated (Fig. 6a), and the roughness of the potential is ultimately responsible for much of the anomalous effects that one observes in these setups. On the other hand, in Eq. (77), anomalous effects can be caused by a nonlinear, drag-velocity-dependent force $\vec{F}(\vec{v})$ in molecular gases. They can also be caused by a shear force and/or the inelasticity of collisions in granular gases (expressed in $\mathcal{J}[\vec{v},\vec{w}|f(t)]$) that produces a non-Gaussian distribution function that couples the kinetic temperature with higher-order moments.
@@
To keep the derivation simple, we analyze a relatively easy case, where there is no deterministic force, and the energy lost in each collision is compensated only by the stochastic homogeneous volume force. This simplification gives a Boltzmann-Fokker-Planck equation of the form
@@
which roughly accounts for the accumulated number of collisions per particle. In the case of thermostatted systems, time can be rescaled with the stationary temperature $T_{s}$,where $T_{s}$ corresponds to the temperature the system reaches when dissipation and injection of energy are balanced; see Sec. 4.2 and 4.3. Dimensionless velocities then follow
@@
$\phi(\vec{c},\vec{j};\tau)\equiv\frac{1}{n}\left[\frac{4T_{t}(t)T_{r}(t)}{mI}\right]^{3/2}f(\vec{v},{\vec{w}};t).$ (86)
@@
Our goal is to use dimensionless variables and write evolution equations for the various temperatures defined above, to characterize crossings and hence anomalous relaxation effects. To achieve this goal, we formally expand the exact solution to the Boltzmann-Fokker-Planck equation Eq. (78) around a Maxwellian distribution, which here is a Gaussian distribution with variances $T_{t}$ and $T_{r}$ for the linear and angular velocities, respectively [156]. Explicitly, we write
$\phi_{M}(c,j)=n\pi^{-3}e^{-c^{2}-j^{2}}.$ (88)
@@
are a dimensionless measure of noise intensity and the temperature associated to a given noise intensity respectively.
Given the temporal evolution of the temperature, we can define various anomalous relaxation phenomena: If, in two relaxation processes with different initial conditions but identical steady state, the corresponding curves of $T(\tau)$ cross, then there is a direct/inverse Mpemba effect. The specific type of the effect depends on the relaxation process – heating or cooling. The Mpemba effect has been studied using numerical, analytical techniques and two simulation methods detailed in the numerical observation; see Sec. 7.3.
@@
where the asterisks ($\vec{v}^{*},\vec{w}^{*}$) indicate precollisional velocities, namely, the velocities that after collision give as results ($\vec{v},\vec{w}$). Note that in the case of molecular gases, the collision kernel is the same, but $\alpha$ must be replaced by $1$ and $\beta$ by ±$\pm 1$, which corresponds to the case of elastic collisions.